OpenStudy (anonymous):

Which proportion has a solution of 8? ten over fifteen equals n over twelve ten over fifteen equals twelve over n ten over n equals twelve over fifteen ten over twelve equals n over fifteen

OpenStudy (anonymous):

first

OpenStudy (anonymous):

a

OpenStudy (anonymous):

10/15=n/12 2/3=n/12 (2/3)*12=n 44*2=n n=8